# Consolidating TinyML Lifecycle with Large Language Models: Reality, Illusion, or Opportunity?

### v1

@@ I Introduction
The rapid development of Internet of Things (IoT) technologies and applications has generated a fast-growing need for the development and deployment of machine learning (ML) models at the edge on resource-constrained devices, a paradigm typically called TinyML. TinyML aims to provide intelligent capabilities to small-size factor and power-efficient devices, unlocking ML-powered applications in smart home automation, health, and industrial monitoring. However, the lifecycle management of TinyML models presents unique challenges that demand high human intervention most of the time.
@@
Our work takes an initial step toward addressing this gap by leveraging LLMs not only for code generation but also to streamline and automate key TinyML lifecycle components, representing an additional contribution to the intersection of Generative AI and IoT. Specifically, we developed a framework that incorporates the natural language understanding and code generation capabilities of GPT-4o, a leading LLM, as a component within a broader system for automating TinyML lifecycle stages. The framework itself plays a critical role in coordinating these stages within the pipeline, managing interactions with the LLM, and adapting each stage’s outputs to ensure compatibility with constrained devices.
@@ II Navigating the TinyML Lifecycle
These challenges highlight the need to rethink various stages of the TinyML lifecycle to enable the effective deployment of ML models on resource-constrained devices like microcontrollers [5]. Fig. 1 compares the traditional ML lifecycle with a generic TinyML lifecycle. Both lifecycles consist of three stages: data collection, development, and deployment & monitoring. However, the TinyML lifecycle involves additional steps and distinctions. For example, TinyML data is typically collected via edge device peripherals, such as sensors, from real or simulated production environments. During the ML model development stage, techniques such as quantization and model compression are applied to address the challenges posed by hardware heterogeneity and constraints specific to each deployment environment. Before deployment, models must be converted to formats compatible with TinyML frameworks (e.g., tflite for TensorFlow Lite [^2] https://www.tensorflow.org/lite/) and compiled into forms that can run on microcontrollers, often through the use of dedicated ML software libraries suitable for embedded ML, requiring C/C++ code or a lightweight runtime to interpret the model on the target device [6].
@@ III The Rise of Language Models
Built upon transformer architectures with self-attention mechanisms, LLMs owe their versatility across various tasks—including text generation, translation, and software development—to their enormous scale and the diversity of data on which they are trained. For example, OpenAI’s GPT-4o model [^3] https://platform.openai.com/docs/models/gpt-4o, which boasts approximately 1.8 trillion parameters across 120 layers, draws on extensive training datasets spanning diverse domains, allowing it to perform complex tasks and adapt to a wide range of applications [7].
@@
The Role of Prompt Engineering. To fully unlock their potential, LLMs often rely on prompt engineering—a technique designed to elicit precise and reliable outputs from these models. Crafting effective prompts enables LLMs to follow structured instructions, producing desired outcomes across a range of complex tasks. Effective prompts typically include role definition, instructions, context, input data, and output indicator. Core prompt engineering techniques, such as Few-Shot Prompting, Chain-of-Thought reasoning, and Self-Consistency prompting, enhance the accuracy and control of LLM responses, especially for multi-step processes [10]. These approaches help to improve LLM performance in tasks like code generation but also address domain-specific challenges by providing clear, step-by-step guidance.
@@
  • Tool Integration and Orchestration: The TinyML lifecycle requires close integration of the LLM with specialized embedded ML libraries, device-specific compilers, and lightweight runtimes. Effective orchestration between these components and the LLM is essential for managing workflows, coordinating data inputs, and ensuring that each output is compatible with the constrained target environment. Middleware LLMs frameworks, such as LangChain [^4] https://www.langchain.com/ for prompt management and LangSmith [^5] https://www.langchain.com/langsmith for monitoring LLM interactions, play a pivotal role in enabling these complex, multi-component interactions.
@@ V Blueprint for Automation
Following the requirements outlined in the previous sections, we introduce our framework designed to integrate OpenAI’s GPT family of LLMs with TinyML tools, aimed at automating key stages of the TinyML lifecycle. We present the framework from two perspectives—the what and the how—emphasizing both its structural components and operational workflow.
@@
The ”how” — Process Workflow. Illustrated in Fig. 3, it focuses on the structured workflow that drives TinyML applications from user input to deployment. The Workflow Stages (left panel) outline the main functional steps: Data Processing, Model Conversion, and Sketch Generation, among others. These stages are orchestrated by the Lifecycle Middleware (detailed in Fig. 2) that bridges user-defined goals with LLM-powered code generation. Meanwhile, the Workflow Logic (right panel) presents a five-step logical flow—from user input acquisition to iterative refinement—ensuring each stage receives feedback for continuous improvement and precision.
At the heart of the Fig. 2 framework is the Lifecycle Middleware, which integrates multiple specialized components. Custom code modules implement the core logic of the framework and facilitate user interactions, while ML software libraries prepare machine learning models for deployment by managing tasks such as conversion and quantization. Dedicated components for LLM integration coordinate interactions with OpenAI’s GPT models, and utility libraries handle essential data processing and manipulation tasks to ensure smooth workflow transitions.
@@
The Custom Code Modules act as the central orchestrator, guiding the flow of user-defined goals—such as requirements, preferences, and specifications—into the LLM environment. It facilitates seamless communication between the LLM and the framework, ensuring that outputs are continuously refined and adapted to align with the hardware and software constraints of TinyML devices. These modules enable the framework to generate outputs that are then processed by specialized embedded ML libraries, such as TFLite, for model conversion and quantization, ensuring alignment with the hardware and software constraints of the target devices. This is achieved by coordinating tasks and managing the iterative exchange with the LLM.
@@
Ultimately, the framework generates essential artifacts—processed datasets, optimized models, and executable sketches—that are fully compatible with TinyML devices.
@@
In each lifecycle stage automated by the framework (Fig. 3 left panel), the workflows follow a consistent logic, though each stage—such as data processing, model conversion, or sketch generation—has specific operations and techniques. By adhering to a standardized workflow (Fig. 3 right panel), each stage begins by gathering stage-specific information. For example, in data processing, the system collects information about the dataset and model purpose, while in model conversion, it gathers details like dataset overview and quantization requirements. Using the gathered input, the framework constructs tailored prompts specific to each stage. These prompts are then passed to the LLM to generate configuration settings or code snippets, which the framework subsequently executes or compiles locally. For instance, the system may execute Python code or compile an Arduino sketch, depending on the stage requirements. To illustrate, in the data processing stage, GPT-4o assists in automating preprocessing steps, including data cleaning, normalization, and augmentation, ensuring that the data is prepared for model training with minimal manual intervention. For model conversion (optimization and quantization), the framework provides scripts and configurations to convert models into formats suitable for deployment on resource-constrained devices, such as TensorFlow Lite. Finally, in the deployment stage (sketch generation), the LLM generates scripts tailored to the specific hardware and application requirements, including code for deploying models on devices like the Arduino Nano 33 BLE. It is worth mentioning that our work focuses specifically on representative stages of the ML lifecycle that are unique to TinyML workflows due to the computing constraints and requirements of resource-constrained devices. These include stages like model optimization and deployment, which necessitate additional operations compared to a traditional ML lifecycle. We deliberately omit other stages, such as model training, as these remain largely identical in both TinyML and traditional ML workflows. For each lifecycle stage, if an error arises during local execution, the system triggers an iterative retry mechanism that engages the LLM with specialized error-handling prompts. This iterative process continues, refining the output until successful execution is achieved or until the maximum retry threshold is reached. The process concludes either with the generation of final artifacts—such as processed datasets, converted models, or Arduino sketches—or with termination if the iteration limit is exceeded.
@@
Fig. 4 shows the interactions between the user, framework, LLM, device compiler, and target device during the deployment stage, including the retry mechanisms that ensure robust code generation and deployment. Instead of displaying the entire prompt code, the snippet on the right provides a structured template summarizing the core components of the prompt, including context, objectives, task-specific instructions, and error handling. This template guides the LLM in generating code that aligns with the hardware constraints and application requirements.
@@ VI Case Study: From Fruit to Functionality
To explore these questions, we conducted a detailed case study involving a fruit classification task on a commonly used IoT device with embedded ML capabilities: the Arduino Nano 33 BLE board. From the ML model perspective, we employed a simple computer vision approach using a convolutional neural network (CNN) model for fruit classification. This model is designed to classify different types of fruits (e.g., apples, bananas, oranges) based on their color variations, leveraging the device’s built-in RGB color sensor.
@@
Empirical Evaluation. We conducted thirty full runs for each lifecycle stage under evaluation—data processing, 8-bit model quantization and conversion, and sketch generation—using GPT-4o (GPT-4o-2024-08-06). Tokens, the smallest units of text input processed by LLMs, served as a key metric for evaluating resource consumption. To ensure controlled testing, the maximum number of iterations for resolving errors through the framework was capped at five.
@@ VII Challenges and Horizons
Enhancing Error Resilience. One significant challenge lies in the system’s limited ability to self-correct, especially in multi-step processes where errors in early stages propagate and compound. A promising direction involves integrating multi-agent LLM systems, combining smaller, locally deployed LLMs with more capable, cloud-based models. Local LLMs could handle error detection and validation, ensuring outputs adhere to expected formats before engaging larger models for more complex tasks. This approach balances privacy, efficiency, and cost, as local models reduce reliance on costly cloud resources. However, computational constraints on TinyML devices make deploying sophisticated models locally challenging, necessitating further optimization and adaptation.
Streamlining Resource-Intensive Processes. Sketch generation emerged as a bottleneck due to high resource consumption and lower success rates. Enhancing prompt engineering and workflow design could address this, focusing on precise, context-aware prompts and iterative feedback loops to reduce errors. Exploring domain-specific languages (DSLs) tailored for target platforms, such as Arduino, could streamline code generation. Hybrid approaches that combine LLM-generated drafts with traditional compilers and manual developer refinement may further improve reliability and efficiency, especially for complex or resource-intensive tasks.
Adaptive Task Allocation for Cost Efficiency. Monetary costs associated with API calls become significant at scale, particularly in dynamic IoT deployments. Adaptive LLM routing offers a solution, directing simpler tasks to smaller, cost-effective models while reserving more complex queries for larger models. This strategy can extend to specialized LLMs for TinyML, enabling dynamic updates or functionality shifts, such as adding new features or reconfiguring device applications. By acting as a capability discovery system, the framework can negotiate with specialized LLMs to identify the most efficient solution, optimizing cost and performance while maintaining adaptability for evolving TinyML requirements.
@@ VIII Conclusions: Reality, Illusion, or Opportunity?
In this paper, we explored the integration of LLMs into the TinyML lifecycle, investigating their potential to automate and streamline key stages through the development of a lifecycle middleware framework tested through a practical IoT development scenario involving a classification use case. Our findings reveal both the promise and challenges of this approach. On the positive side, LLMs demonstrate the ability to improve development efficiency, as seen in tasks like model quantization and data preprocessing. However, limitations remain: reliability issues in complex tasks such as sketch generation, constrained generalizability across hardware platforms, and significant resource demands in terms of computation and cost. These challenges highlight the need for further refinement, including enhanced prompt engineering, specialized model fine-tuning, and the development of advanced symbolic reasoning mechanisms to improve error correction, iterative refinement, and reasoning capabilities. Despite these hurdles, the opportunity for transformation is clear. By addressing these challenges and expanding the framework to include additional lifecycle stages and more diverse hardware, LLMs can become a cornerstone of TinyML automation. This work lays the groundwork for leveraging LLMs alongside traditional tools, paving the way for more efficient, scalable, and accessible embedded IoT ML workflows.

### v2

@@ I Introduction
The rapid development of Internet of Things (IoT) technologies and applications has generated a fast-growing need for the development and deployment of machine learning (ML) models at the edge on resource-constrained devices, a paradigm typically called TinyML [1]. TinyML refers to the deployment of machine learning models on resource-constrained devices such as microcontrollers. Unlike traditional ML, which typically runs on powerful cloud or edge servers, TinyML models must operate under tight limitations in memory, compute power, and energy, making their development and lifecycle management significantly more complex [2]. TinyML aims to provide intelligent capabilities to small-size factor and power-efficient devices, unlocking ML-powered applications in smart home automation, health, and industrial monitoring. However, the lifecycle management of TinyML models presents unique challenges that demand high human intervention most of the time.
@@
Our work takes an initial step toward addressing this gap by leveraging LLMs not only for code generation but also to streamline and automate key TinyML lifecycle components, representing an additional contribution to the intersection of Generative AI and IoT. Specifically, we developed a framework that incorporates the natural language understanding and code generation capabilities of GPT-4o, a leading LLM, as a component within a broader system for automating TinyML lifecycle stages. The framework itself plays a critical role in coordinating these stages within the pipeline, managing interactions with the LLM, and adapting each stage’s outputs to ensure compatibility with constrained devices. Unlike traditional TinyML toolchains (e.g., TensorFlow Lite [4], Edge Impulse [5]) that require specialized human involvement across multiple steps, our framework focuses on minimizing developer effort by using LLMs to automate key tasks such as configuration generation, conversion scripting, and deployment sketch creation. Rather than replacing these toolchains, our framework is designed to be complementary, offering an automation layer particularly beneficial in hyperscale IoT deployments, where managing tailored workflows across numerous heterogeneous devices can become increasingly cumbersome.
@@ II Navigating the TinyML Lifecycle
These challenges highlight the need to rethink various stages of the TinyML lifecycle to enable the effective deployment of ML models on resource-constrained devices like microcontrollers [1]. Fig. 1 compares the traditional ML lifecycle with a generic TinyML lifecycle. Both lifecycles consist of three stages: data collection, development, and deployment & monitoring. However, the TinyML lifecycle involves additional steps and distinctions. For example, TinyML data is typically collected via edge device peripherals, such as sensors, from real or simulated production environments. During the ML model development stage, techniques such as quantization and model compression are applied to address the challenges posed by hardware heterogeneity and constraints specific to each deployment environment. Before deployment, models must be converted to formats compatible with TinyML frameworks (e.g., tflite for TensorFlow Lite) and compiled into forms that can run on microcontrollers, often through the use of dedicated ML software libraries suitable for embedded ML, requiring C/C++ code or a lightweight runtime to interpret the model on the target device [6].
@@ III The Rise of Language Models
Built upon transformer architectures with self-attention mechanisms, LLMs owe their versatility across various tasks—including text generation, translation, and software development—to their enormous scale and the diversity of data on which they are trained. For example, OpenAI’s GPT-4o model, which boasts approximately 1.8 trillion parameters across 120 layers, draws on extensive training datasets spanning diverse domains, allowing it to perform complex tasks and adapt to a wide range of applications [7].
@@
The Role of Prompt Engineering. To fully unlock their potential, LLMs often rely on prompt engineering, which is a technique designed to elicit precise and reliable outputs from these models. Crafting effective prompts enables LLMs to follow structured instructions, producing desired outcomes across a range of complex tasks. Effective prompts typically include role definition, instructions, context, input data, and output indicator. Core prompt engineering techniques, such as Few-Shot Prompting, Chain-of-Thought reasoning, and Self-Consistency prompting, enhance the accuracy and control of LLM responses, especially for multi-step processes [10]. These approaches help to improve LLM performance in tasks like code generation but also address domain-specific challenges by providing clear, step-by-step guidance.
@@
  • Tool Integration and Orchestration: The TinyML lifecycle requires close integration of the LLM with specialized embedded ML libraries, device-specific compilers, and lightweight runtimes. Effective orchestration between these components and the LLM is essential for managing workflows, coordinating data inputs, and ensuring that each output is compatible with the constrained target environment. Middleware LLMs frameworks, such as LangChain for prompt management and LangSmith for monitoring LLM interactions, play a pivotal role in enabling these complex, multi-component interactions.
@@ V Blueprint for Automation
Following the requirements outlined in the previous sections, we introduce our framework designed to integrate OpenAI’s GPT family of LLMs with TinyML tools, aimed at automating key stages of the TinyML lifecycle. We present the framework from two perspectives regarding the what and the how, emphasizing both its structural components and operational workflow.
@@
The ”how” — Process Workflow. Illustrated in Fig. 3, it focuses on the structured workflow that drives TinyML applications from user input to deployment. The Workflow Stages (left panel) outline the main functional steps: Data Processing, Model Conversion, and Sketch Generation, among others. These stages are orchestrated by the Lifecycle Middleware (detailed in Fig. 2) that bridges user-defined goals with LLM-powered code generation. Meanwhile, the Workflow Logic (right panel) presents a five-step logical flow from user input acquisition to iterative refinement, ensuring each stage receives feedback for continuous improvement and precision.
At the heart of the Fig. 2 framework is the Lifecycle Middleware, which integrates multiple specialized components. Custom code modules implement the core logic of the framework and facilitate user interactions, while ML software libraries prepare machine learning models for deployment by managing tasks such as conversion and quantization. Dedicated components for LLM integration coordinate interactions with OpenAI’s GPT models, and utility libraries handle essential data processing and manipulation tasks to ensure smooth workflow transitions. Importantly, once the user provides the dataset, the subsequent workflow is fully automated. The framework leverages a set of curated prompt templates, each corresponding to a different lifecycle stage (e.g., data processing, model conversion, deployment). These templates are dynamically adapted based on the task context and feedback signals, such as errors captured during execution, allowing the system to invoke the LLM autonomously and progress through the pipeline without requiring further human intervention.
@@
The Custom Code Modules act as the central orchestrator, guiding the flow of user-defined goals such as requirements, preferences, and specifications into the LLM environment. It facilitates seamless communication between the LLM and the framework, ensuring that outputs are continuously refined and adapted to align with the hardware and software constraints of TinyML devices. These modules enable the framework to generate outputs that are then processed by specialized embedded ML libraries, such as TFLite, for model conversion and quantization, ensuring alignment with the hardware and software constraints of the target devices. This is achieved by coordinating tasks and managing the iterative exchange with the LLM.
@@
Ultimately, the framework generates essential artifacts, including processed datasets, optimized models, and executable sketches that are fully compatible with TinyML devices.
@@
In each lifecycle stage automated by the framework (Fig. 3 left panel), the workflows follow a consistent logic, though each stage, such as data processing, model conversion, or sketch generation, has specific operations and techniques. By adhering to a standardized workflow (Fig. 3 right panel), each stage begins by gathering stage-specific information. For example, in data processing, the system collects information about the dataset and model purpose, while in model conversion, it gathers details like dataset overview and quantization requirements. Using the gathered input, the framework constructs tailored prompts specific to each stage. These prompts are then passed to the LLM to generate configuration settings or code snippets, which the framework subsequently executes or compiles locally. For instance, the system may execute Python code or compile an Arduino sketch, depending on the stage requirements. To illustrate, in the data processing stage, GPT-4o assists in automating preprocessing steps, including data cleaning, normalization, and augmentation, ensuring that the data is prepared for model training with minimal manual intervention. For model conversion (optimization and quantization), the framework provides scripts and configurations to convert models into formats suitable for deployment on resource-constrained devices, such as TensorFlow Lite. Finally, in the deployment stage (sketch generation), the LLM generates scripts tailored to the specific hardware and application requirements, including code for deploying models on devices like the Arduino Nano 33 BLE. It is worth mentioning that our work focuses specifically on representative stages of the ML lifecycle that are unique to TinyML workflows due to the computing constraints and requirements of resource-constrained devices. These include stages like model optimization and deployment, which necessitate additional operations compared to a traditional ML lifecycle. We deliberately omit other stages, such as model training, as these remain largely identical in both TinyML and traditional ML workflows. For each lifecycle stage, if an error arises during local execution, the system triggers an iterative retry mechanism that engages the LLM with specialized error-handling prompts. This iterative process continues, refining the output until successful execution is achieved or until the maximum retry threshold is reached. The process concludes either with the generation of final artifacts, such as processed datasets, converted models, or Arduino sketches, or with termination if the iteration limit is exceeded.
@@
Fig. 4 shows the interactions between the user, framework, LLM, device compiler, and target device during the sketch deployment stage, including the retry mechanisms that ensure robust code generation and deployment. It is worth emphasizing that the framework includes a feedback-driven retry mechanism: when an operation fails (e.g., due to compilation or deployment errors, the system parses the error and updates the prompt accordingly. This process is handled automatically, enabling the LLM to regenerate code that better aligns with the hardware constraints or resolves the encountered issue—without manual prompt rewriting. Instead of displaying the entire prompt code, the snippet on the right provides a structured template summarizing the core components of the prompt, including context, objectives, task-specific instructions, and error handling. This template guides the LLM in generating code that aligns with the hardware constraints and application requirements.
@@ VI Case Study: From Fruit to Functionality
To explore these questions, we conducted a detailed case study involving a fruit classification task—representative of computer vision workloads—on the Arduino Nano 33 BLE board. This use case was chosen because it aligns with an officially supported Arduino example, allowing us to validate our LLM-generated sketch against a reliable reference implementation. While simple in nature, it includes the typical lifecycle steps (data processing, quantization, and hardware-specific code generation), and our framework is readily extensible to more complex datasets or embedded ML applications. From the ML model perspective, we employed a simple computer vision approach using a convolutional neural network (CNN) model for fruit classification. This model is designed to classify different types of fruits (e.g., apples, bananas, oranges) based on their color variations, leveraging the device’s built-in RGB color sensor.
@@
Empirical Evaluation. We conducted thirty full runs for each lifecycle stage under evaluation—data processing, 8-bit model quantization and conversion, and sketch generation—using GPT-4o (GPT-4o-2024-08-06). Tokens, the smallest units of text input processed by LLMs, served as a key metric for evaluating resource consumption. To ensure controlled testing, the maximum number of iterations for resolving errors through the framework was capped at five (customizable), preventing excessive retries while allowing prompt adaptation to converge within practical bounds. A stage was considered successful if it completed without LLM failure or compilation error within the retry limit. Execution time was measured from the initial LLM query to the successful completion of the stage. Per-stage results were aggregated over 30 runs, and we reported mean, range, and success rate for each stage to reflect both central tendency and variability. The implementation of the framework spans approximately 1,578 lines of Python code, covering the automation logic, prompt handling, error parsing, and interaction with the LLM API and device toolchain (i.e., Arduino Command Line Tool in this case). It is worth highlighting, this end-to-end process introduces no additional computational burden on the target device, as all LLM interactions and code generation occur off-device, and the resulting deployment artifacts are nearly identical to those produced via standard TinyML workflows.
@@ VII Challenges and Horizons
Enhancing Error Resilience. One significant challenge lies in the system’s limited ability to self-correct, especially in multi-step processes where errors in early stages propagate and compound. A promising direction involves considering LLM-based multi-agent systems, combining smaller, locally deployed LLMs with more capable, cloud-based models. Local LLMs could handle error detection and validation, ensuring outputs adhere to expected formats before engaging larger models for more complex tasks. Although performance–cost–bandwidth trade-offs should be evaluated on a case-by-case basis, this approach can help balance privacy, efficiency, and cost, as local models reduce reliance on costly cloud resources. However, computational constraints on TinyML devices make deploying sophisticated models locally challenging, necessitating further optimization and adaptation. In such cases, an alternative strategy could involve equipping the framework, which effectively stands between the end device and the cloud, with a lightweight, embedded LLM acting as a local proxy. This would enable more responsive coordination and selective filtering, improving latency and efficiency in distributed inference workflows. In addition, the use of adaptive prompt templates, where prompt structure evolves dynamically based on execution outcomes and feedback, may further enhance the system’s ability to self-correct. Such templates can prevent repeated failures by refining LLM instructions contextually, improving reliability across iterations.
Streamlining Resource-Intensive Processes. Sketch generation emerged as a bottleneck due to high resource consumption and lower success rates. Enhancing prompt engineering and workflow design could address this, focusing on precise, context-aware prompts and iterative feedback loops to reduce errors. As mentioned above, adaptive prompt templates represent a promising technique in this direction, enabling more reliable and tailored code generation through dynamic prompt evolution. Additionally, exploring domain-specific languages (DSLs) tailored for target platforms, such as Arduino, could streamline code generation. As an example, DSLs inspired by symbolic programming languages, such as Prolog, have been investigated for improving reasoning and control logic in constrained environments, offering an avenue to complement LLM-based generation with more structured and explainable constructs. Finally, hybrid approaches that combine LLM-generated drafts with traditional compilers and manual developer refinement may further improve reliability and efficiency, especially for complex or resource-intensive tasks.
Adaptive Task Allocation for Cost Efficiency. Monetary costs associated with API calls become significant at scale, particularly in dynamic IoT deployments. Adaptive LLM routing offers a solution, directing simpler tasks to smaller, cost-effective models while reserving more complex queries for larger models. For instance, tasks such as data processing or model conversion—which in our evaluation showed higher success rates—can be routed to LLMs with lower-cost APIs. This introduces a trade-off: slightly lower reliability may require more iterations, but can yield meaningful cost savings. This strategy can extend to specialized LLMs for TinyML, enabling dynamic updates or functionality shifts, such as adding new features or reconfiguring device applications. For example, if an application initially performs object detection but later requires object tracking, the framework can route the request to a model more specialized in temporal or visual sequence tasks. By acting as a capability discovery system, the framework can negotiate with specialized LLMs to identify the most efficient solution, optimizing cost and performance while maintaining adaptability for evolving TinyML requirements.
@@ VIII Conclusions: Reality, Illusion, or Opportunity?
In this paper, we explored the integration of LLMs into the TinyML lifecycle, investigating their potential to automate and streamline key stages through the development of a lifecycle middleware framework tested through a practical IoT development scenario involving a classification use case. Our findings reveal both the promise and challenges of this approach. On the positive side, LLMs demonstrate the ability to improve development efficiency, as seen in tasks like model quantization and data preprocessing. However, limitations remain: reliability issues in complex tasks such as sketch generation, constrained generalizability across hardware platforms, and significant resource demands in terms of computation and cost. These challenges highlight the need for further refinement, including enhanced prompt engineering, specialized model fine-tuning, and the development of advanced symbolic reasoning mechanisms to improve error correction, iterative refinement, and reasoning capabilities. Despite these hurdles, the opportunity for transformation is clear. By addressing these challenges and expanding the framework to include additional lifecycle stages and more diverse hardware, LLMs can become a cornerstone of TinyML automation. This work lays the groundwork for leveraging LLMs alongside traditional tools, paving the way for more efficient, scalable, and accessible embedded IoT ML workflows. Through the development of this end-to-end framework for automating core TinyML lifecycle stages using LLMs, we contribute a novel methodology that reduces developer effort and highlights the feasibility of language model–driven embedded ML workflows. Looking forward, open questions remain about the long-term generalizability of LLM-generated code across highly heterogeneous hardware platforms, and how such models can reliably adapt to evolving application needs with minimal supervision. Addressing these questions may define the next phase of LLM-enabled intelligence at the edge.